# Betelgeuse, Betelgeuse, Betelgeuse, Betel-buddy? Constraints on the dynamical companion to  Orionis from HST

### v1

@@ 4.4 Implications for the Companion
Based on our comparison of observations of the Betelgeuse system during this campaign to characteristic spectral features from YSOs with mass determinations in the ULLYSES database, we confidently exclude a mass of for a YSO companion. We do not definitively exclude a mass of , but find the likelihood to be marginal. Masses below are not excluded. Figure 9 illustrates these scenarios with stellar evolutionary tracks (left panel) from MESA Isocrones and Stellar Tracks (MIST; Choi et al., 2016; Dotter, 2016; Paxton et al., 2011, 2013, 2015, 2018, left panel) at the masses given for the ULLYSES objects shown in Figure 7. The tracks shown span the Hayashi track through truncation at twice Betelgeuseâs approximate age, 20 Myr (prior to the zero-age main sequence in most cases). Lines of constant age (isochrones, red) at 5 Myr and 15 Myr are shown, with thick curves between (purple-blue-green) liberally indicating approximate plausible ages of . The region bounded by the red isochrones and purple-blue-green evolutionary tracks thus approximately illustrates the marginal (dashed) or fully consistent (solid) remaining locus of the companion (with redder and fainter companions than those shown also fully consistent).
@@
The mass resolution of the ULLYSES observations we compare to, and more importantly, the extreme diversity of YSO luminosity in UV and high-energy emission (Preibisch et al. 2005; Preibisch & Feigelson 2005; Getman et al. 2005 and seen in our Figure 1), make it difficult to place strong constraints on the companion mass. Nevertheless, it is interesting that we only encounter true exclusion for a relatively high mass of (dotted green curve), and marginal constraints for the 1.1â1.5 cases (dashed blue-green curves). The fact that a YSO as massive as is not fully excluded leaves open the need for future resolved and pan-chromatic observing campaigns to place more detailed constraints on the companionâs physical nature and interaction with its environment, or, to identify a different cause of the Long Secondary Period persistent in Betelgeuseâs photometry, astrometry, and radial velocity from which the putative companion was inferred.
@@
Likewise, flux-calibrated libraries like XSL (Chen et al., 2014; Manara et al., 2017) extended to the FUV would be valuable for constraining quiet or near-MS companions in the range which arenât well-captured in the ULLYSES data set, as broader libraries across a wider parameter space and absolute flux calibration could facilitate a more realistic detection threshold. Either way, the most promising observational path forward for direct detection remains long-baseline observations and analysis of historical observations of the resolved stellar surface and circumstellar envelope, specifically informed by timing relative to the LSP phase. Likewise, future and proposed FUV missions such as the Large UV/Optical/IR Surveyor (LUVOIR, The LUVOIR Team 2019) and the Cosmological Advanced Survey Telescope for Optical and ultraviolet Research (CASTOR, CĂŽte et al. 2012) may provide additional, deeper constraints, in particular on the UV continuum emission of Betelgeuse as well as higher-contrast features from the companion at favorable orbital configurations.
@@ 5 Conclusions
In this work, we detailed our HST STIS E140M observations of Betelgeuse during mid-to-late November 2024, near its recently-predicted (Goldberg et al., 2024; MacLeod et al., 2025) companionâs receding limb passage. This resulted in a non-detection: we do not find any statistically significant line emission in the FUV which could be identified as originating from the companion at the expected RV offset from Betelgeuse (km/s redshift, corresponding to a AU orbit with a d period), within flux limits of ergâs-1âcm-2âĂ -1 which increase with increasing wavelength (across the full band, the average flux is ergâs-1âcm-2âĂ -1, and the error is ergâs-1âcm-2âĂ -1). There is likewise no statistically significant contribution to continuum emission from a companion (i.e. the continuum differences between quadrants are consistent with HSTâs error), and continuum-like emission from Betelgeuse is just barely at the level compared to the error. Comparing to a set of YSO spectra taken with STIS in the ULLYSES database, we determine that higher-mass scenarios for the companion are ruled out, placing a spectroscopic mass constraint of and favoring masses , with acknowledgment of the great diversity in YSO spectra even within a given mass or spectral-type bin. This remains consistent with the lower-mass end of estimates fit from the STELLA RV only (Goldberg et al., 2024), but favors masses in the range as predicted by the long-baseline RV analysis (MacLeod et al., 2025).
@@
Although we cannot uniquely identify spectroscopic features from the companion itself, we see slight differences in the spectra for each quadrant of our HST observations which loosely correspond on the sky to the Northern, Southern, Eastern, and Western quadrants of Betelgeuseâs chromosphere and circumstellar envelope. Though these differences are only slightly above the background, they likely indicate spatial variation across the chromosphere and circumstellar envelope. This is expected given the large-scale convective nature of RSG envelopes (see, e.g., recent review by Chiavassa et al., 2024), and is consistent with whatâs seen in interferometric studies of Betelgeuse (e.g. Harper et al., 2001; OâGorman et al., 2015; MontargĂšs et al., 2014, 2021; Kervella et al., 2009, 2018; Haubois et al., 2009, 2019, 2023; Kravchenko et al., 2021) as well as other work on spatially resolved UV observations of Betelgeuse and its chromosphere (e.g. Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Harper & Brown, 2006; Dupree et al., 2005, 2020; Matthews & Dupree, 2022). These differences are especially visible in the Ly- line (1215.67Ă ), though this feature is convolved with intense geocoronal emission and we were not able to cleanly separate the stellar Ly- profile. While the line itself is broadened by 10s of km/s, the velocity offsets of the Ly- line peaks in the four quadrants are consistent with the smeared-and-projected km/s rotational velocity (e.g. Dupree et al., 1987; Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Kervella et al., 2009, 2018). These spatial RV variations are also seen in the other emission features. Further modeling and future spatially-resolved observations may connect spectroscopic differences to spatial differences in chromospheric emission from Betelgeuse.
@@
Additional X-ray observations with Chandra, presented in OâGrady et al. (2025), exclude a compact object (white dwarf or neutron star), and constrain the X-ray emission coming from a possible YSO companion with larger uncertainties arising from uncertainty in the X-ray absorption along the line-of-sight as well as the multiple-orders-of-magnitude differences in X-ray emission from YSOs even of the same spectral class. Therefore, future observational campaigns remain necessary for direct detection of , in order to confirm its presence and place constraints on its spectroscopic nature, interaction with the stellar wind, and any possible connection to dust formation such as during Great Dimming-like episodes. Finally, we re-emphasize the importance of considering the stellar variability cycles, and in particular Long Secondary Periods, when constructing spectroscopic campaigns for Betelgeuse and for all other binary and variable stars.
@@ Acknowledgements
We thank Morgan MacLeod for invaluable insight and discussions. We likewise acknowledge helpful conversations with Andrea Antoni, Matteo Cantiello, Thavisha Dharmawardena, Zarina Dhillon, Kareem El-Badry, Brian Metzger, Michael Palumbo, Adiv Paradise, Mathieu Renzo, and Lieke van Son. J.A.G. is supported by a Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. A.O. and B.O. gratefully acknowledge support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. M.J. gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commissionâs Widening Fellowship. This project has received funding from the European Unionâs Horizon 2020 research and innovation programme. This research was supported by the âSeismoLabâ KKP-137523 Ălvonal grant and by the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH). M.R.D. acknowledges support from the NSERC through grant RGPIN-2019-06186, the Ontario Ministry of Colleges and Universities through grant ER22-17-164, the Canada Research Chairs Program, and the Dunlap Institute at the University of Toronto.
@@
We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). This research made use of NASAâs Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.

### v2

@@ 4.3 Implications for the Companion
Based on our comparison of observations of the Betelgeuse system during this campaign to characteristic spectral features from YSOs with mass determinations in the ULLYSES database, we confidently exclude a mass of for a YSO companion. We do not definitively exclude a mass of , but find the likelihood to be marginal. Masses below are not excluded. Figure 9 illustrates these scenarios with stellar evolutionary tracks (left panel) from MESA Isocrones and Stellar Tracks (MIST; Choi et al., 2016; Dotter, 2016; Paxton et al., 2011, 2013, 2015, 2018, left panel) at the masses given for the ULLYSES objects shown in Figure 8. The tracks shown span the Hayashi track through truncation at twice Betelgeuse’s approximate age, 20 Myr (prior to the zero-age main sequence in most cases). Lines of constant age (isochrones, red) at 5 Myr and 15 Myr are shown, with thick curves between (purple-blue-green) liberally indicating approximate plausible ages of . The region bounded by the red isochrones and purple-blue-green evolutionary tracks thus approximately illustrates the marginal (dashed) or fully consistent (solid) remaining locus of the companion (with redder and fainter companions than those shown also fully consistent).
@@
The mass resolution of the ULLYSES observations we compare to, and more importantly, the extreme diversity of YSO luminosity in UV and high-energy emission (Preibisch et al. 2005; Preibisch & Feigelson 2005; Getman et al. 2005 and seen in our Figure 1), make it difficult to place strong constraints on the companion mass. Nevertheless, it is interesting that we only encounter true exclusion for a relatively high mass of (dotted green curve), and marginal constraints for the 1.1–1.5 cases (dashed blue-green curves). The fact that a YSO as massive as is not fully excluded leaves open the need for future resolved and pan-chromatic observing campaigns to place more detailed constraints on the companion’s physical nature and interaction with its environment, or, to identify a different cause of the Long Secondary Period persistent in Betelgeuse’s photometry, astrometry, and radial velocity from which the putative companion was inferred.
@@
Likewise, flux-calibrated libraries like XSL (Chen et al., 2014; Manara et al., 2017) extended to the FUV would be valuable for constraining quiet or near-MS companions in the range which aren’t well-captured in the ULLYSES data set, as broader libraries across a wider parameter space and absolute flux calibration could facilitate a more realistic detection threshold. Either way, a promising observational path towards tighter constraints on the companion’s mass and surface proerties remains long-baseline observations and analysis of historical observations of the resolved stellar surface and circumstellar envelope, specifically informed by timing relative to the LSP phase. Likewise, future and proposed FUV missions such as the Large UV/Optical/IR Surveyor (LUVOIR, The LUVOIR Team 2019) and the Cosmological Advanced Survey Telescope for Optical and ultraviolet Research (CASTOR, Côte et al. 2012) may provide additional, deeper constraints, in particular on the UV continuum emission of Betelgeuse as well as higher-contrast features from the companion at favorable orbital configurations.
While this work was in the refereeing process, Howell et al. (2025) reported the probable claimed detection, via speckle imaging, of a source 6 magnitudes () fainter than Betelgeuse in the B band, with location and timing predicted by Goldberg et al. (2024) and MacLeod et al. (2025).121212Sadly, the measurement errors in HST (seen in Figure 7 and the right panel of Figure 9) are far too large to observe a differences in UV continuum flux across quadrants. The authors identify the companion as an young stellar object approaching but not yet on its Main Sequence.
This scenario is prima facie at the margin of the range of YSO masses which we can exclude on the basis of a spectroscopic non-detection. However, chromospheric activity typically decreases as stars age closer to the Main Sequence (Preibisch & Feigelson, 2005; Linsky, 2017). Additionally, for a YSO companion interacting with Betelgeuse, spin-down processes may reduce the spin-controlled magnetic activity to below that of a typical pre-MS star. This may impact the emission limits on possible companion types. For example, since we expect a YSO companion to be stripped of any residual natal disk as discussed in §2.3, that may reduce any disk-locking that would otherwise spin up the companion prior to the disk-free T-Tauri stage. Similarly, tidal coupling between Betelgeuse and as discussed by MacLeod et al. (2025) would reduce the spin of the companion, as Betelgeuse’s spin period is slower than the companion’s orbital period. These scenarios both may suppress the companion’s spin and thus expected UV emission.
Thus, the specific scenario discussed by Howell et al. (2025) is indeed cleanly within the constraints placed of our observations with HST. We therefore eagerly await upcoming efforts towards direct detection with speckle imaging, adaptive optics, and other high-resolution high-contrast observational techniques, in order to confirm the anticipated non-detection at occultation and detection at approaching quadrature.
@@ 5 Conclusions
In this work, we detailed our HST STIS E140M observations of Betelgeuse during mid-to-late November 2024, near its recently-predicted (Goldberg et al., 2024; MacLeod et al., 2025) companion’s receding limb passage. This resulted in a non-detection: we do not find any statistically significant line emission in the FUV which could be identified as originating from the companion at the expected RV offset from Betelgeuse (km/s redshift, corresponding to a AU orbit with a d period), within flux limits of erg s-1 cm-2 Å-1 which increase with increasing wavelength (across the full band, the average flux is erg s-1 cm-2 Å-1, and the error is erg s-1 cm-2 Å-1). There is likewise no statistically significant contribution to continuum emission from a companion (i.e. the continuum differences between quadrants are consistent with HST’s error), and continuum-like emission from Betelgeuse is just barely at the level compared to the error. Comparing to a set of YSO spectra taken with STIS in the ULLYSES database, we determine that higher-mass scenarios for the companion are ruled out, placing a spectroscopic mass constraint of and favoring masses , with acknowledgment of the great diversity in YSO spectra even within a given mass or spectral-type bin. This remains consistent with the lower-mass end of estimates fit from the STELLA RV only (Goldberg et al., 2024), but favors masses in the range as predicted by the long-baseline RV analysis (MacLeod et al., 2025).
@@
Although we cannot uniquely identify spectroscopic features from the companion itself, we see slight differences in the spectra for each quadrant of our HST observations which loosely correspond on the sky to the Northern, Southern, Eastern, and Western quadrants of Betelgeuse’s chromosphere and circumstellar envelope. Though these differences are only slightly above the background, they likely indicate spatial variation across the chromosphere and circumstellar envelope. This is expected given the scale of the vigorously convective nature of RSG envelopes (Stothers 1972, e.g.; Goldberg et al. 2022, e.g.; Chiavassa et al. 2009, 2011, 2024, e.g.; Ma et al. 2025, e.g.). These asymmetries are also consistent with resolved interferometric observations of Betelgeuse (e.g. Harper et al., 2001; O’Gorman et al., 2015; Montargès et al., 2014, 2021; Kervella et al., 2009, 2018; Haubois et al., 2009, 2019, 2023; Kravchenko et al., 2021), as well as other work on spatially resolved UV observations of Betelgeuse and its chromosphere (e.g. Gilliland & Dupree, 1996; Uitenbroek et al., 1998; Lobel & Dupree, 2001; Harper & Brown, 2006; Dupree et al., 2005, 2020; Matthews & Dupree, 2022). Further modeling and future spatially-resolved observations may connect spectroscopic differences to spatial differences in chromospheric emission from Betelgeuse.
@@
Additional X-ray observations with Chandra, presented in O’Grady et al. (2025), exclude a compact object (white dwarf or neutron star), and constrain the X-ray emission coming from a possible YSO companion with larger uncertainties arising from uncertainty in the X-ray absorption along the line-of-sight as well as the multiple-orders-of-magnitude differences in X-ray emission from YSOs even of the same spectral class. Therefore, future observational campaigns remain necessary for direct detection of , in order to confirm its presence and place constraints on its spectroscopic nature, interaction with the stellar wind, and any possible connection to dust formation such as during Great Dimming-like episodes. Finally, we re-emphasize the importance of considering the stellar variability cycles, and in particular Long Secondary Periods, when constructing spectroscopic campaigns for Betelgeuse and for all other binary and variable stars.
@@ Acknowledgements
We thank Morgan MacLeod for invaluable discussions. We likewise acknowledge helpful conversations with Andrea Antoni, Matteo Cantiello, Thavisha Dharmawardena, Zarina Dhillon, Kareem El-Badry, Brian Metzger, Michael Palumbo, Adiv Paradise, Mathieu Renzo, and Lieke van Son. We are forever grateful to Bill Paxton for all he has done for the astrophysical community.
@@
J.A.G. is supported by a Flatiron Research Fellowship. The Flatiron Institute is supported by the Simons Foundation. A.O. and B.O. gratefully acknowledge support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. M.J. gratefully acknowledges funding of MATISSE: Measuring Ages Through Isochrones, Seismology, and Stellar Evolution, awarded through the European Commission’s Widening Fellowship. This project has received funding from the European Union’s Horizon 2020 research and innovation programme. This research was supported by the ‘SeismoLab’ KKP-137523 Élvonal grant and by the NKFIH excellence grant TKP2021-NKTA-64 of the Hungarian Research, Development and Innovation Office (NKFIH). M.R.D. acknowledges support from the NSERC through grant RGPIN-2019-06186, the Ontario Ministry of Colleges and Universities through grant ER22-17-164, the Canada Research Chairs Program, and the Dunlap Institute at the University of Toronto.
@@
We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). This research made use of NASA’s Astrophysics Data System Bibliographic Services, the pathfinder framework (Iyer et al., 2024) as well as of the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France.
Appendix A Nondetection of Lyman- Emission from companion
The Ly- (1215.67Å) emission line is another often-shown feature in the FUV (Linsky, 2017; Carpenter et al., 2018; Roman-Duval et al., 2020, 2025). However, this feature is dominated by contamination from geocoronal emission from Earth’s atmosphere.We attempted, unsuccessfully, to separate possible source and geocoronal Ly- emission profiles by extracting cross-dispersion slices of the 2D spectra, including regions centered on the order traces where Ly- appears along with adjacent background regions on either side. Ly- appears in two adjacent orders in our E140M spectra. For one of the orders, the center trace spectrum shows evidence of a roughly 2.5 km s-1 shift relative to the background Ly- spectrum, but no such signal was detected in the other order.
Nonetheless, the companion’s expected radial velocity shift is larger than the approximate width (35 km/s) of the geocoronal Ly- line, so if the companion were strongly emitting in Ly- it could in-principle be visible here. Therefore, for completeness, Figure 10 shows spectra near the Ly- feature from each of the four observed quadrants, along with the 2011 observations by Carpenter et al. (2018) shifted to the RV of Betelgeuse at the time of our 2024 observations. There is no discernable Ly- emission at the expected redshift of the companion.
Due to the geocoronal origin of the strong Ly- features here, the km/s offset between the 2011 Ly- and the median 2024 Ly- can be attributed to differences in velocity about the Earth’s orbit projected toward Betelgeuse between the different epochs. The broad nature of the Ly- features is likewise attributed to the diffuse sky-glow filling the 0.2x0.2 aperture. Interestingly, Ly- is the only emission feature throughout the entire spectrum that displays this large of a velocity offset between the 2024 and 2011 observations; all other features are only offset by 2.5 km/s between the 2011 and our 2024 observations, owing to the RV difference in Betelgeuse between these dates. This includes the 1302-1306Å OI lines, which are also contaminated by geocoronal emission.
We also compared with template spectra (not shown here) of YSO’s from ULYSSES (Roman-Duval et al., 2020, 2025), but those spectra are likewise contaminated by geocoronal emission. Without clean expectations for the uncontaminated Ly- emission profiles of possible companion scenarios, it is unfortunately impossible to place mass constraints from the lack of observed Ly- from the companion.